OpenStudy (anonymous):

Please Help. 1.Write the equation of the line which passes through (5, –2) and is parallel to x = 4. 2.Write the equation of the line which passes through (2, 1) and is perpendicular to x = –2. 3.Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form. 4.Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form. Need to shw work

OpenStudy (anonymous):

nothing to show for the first one \(x=4\) is a vertical line, so you are asked for the equation of another vertical line, this one through the point \((5,-2)\) you write \(x=\text{the first coordinate}\)

OpenStudy (anonymous):

For the first one is it x = 5

OpenStudy (anonymous):

second one is similar, but since \(x=-2\) is a vertical line, you want the equation for a horizontal line horizontal line looks like \(y=\text{some number}\) and since you want it to go through the point \((2,1)\) you write \(y=\text{second coordinate}\)

OpenStudy (anonymous):

yes, first one is \(x=5\)

OpenStudy (anonymous):

you got the second one?

OpenStudy (anonymous):

So y = 1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

third one requires some actual work do you know what the slope of \(y=-x+6\) is ?

OpenStudy (anonymous):

would the slope be a negative number

OpenStudy (anonymous):

yes, in fact it would be \(-1\) since \(y=-x+6\) looks like \(y=-1\times x+6\) general form is \(y=mx+b\) where \(m\) is the slope, so in this case \(m=-1\)

OpenStudy (anonymous):

now we interpret the question as saying "find the equation of the line with slope \(-1\) through the point \((-4,2)\) " do you know how to do that?

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

ok this is what you need \[y-y_1=m(x-x_1)\] this is the "point-slope" formula you have a point, it is \((-4,2)\) this means \(x_1=-4,y_1=2\) and you have the slope \(m=-1\) plug these numbers directly in to the formula can you do that or would you like me to show it?

OpenStudy (anonymous):

is it y - 2 = -1 (x - 4)

OpenStudy (anonymous):

oops, i was wrong it is \(y-2=-1(x-(-4))\) or \[y-2=-(x+4)\]

OpenStudy (anonymous):

I got it

OpenStudy (anonymous):

then you still have one more job to do you have to write it in the form \(y=mx+b\) distribute the minus sign on the right and get \[y-2=-x-4\] then add 2 to both sides to get \[y=-x-2\]

OpenStudy (anonymous):

these steps are always the same, so with a little practice you will be able to do them swiftly

OpenStudy (anonymous):

Thank you, your a big help. I actually understand it more thn reading it in my lesson.

OpenStudy (anonymous):

yw try the next one' question is Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form. slope of \(y=4x+7\) is 4, slope of perpendicular line is the "negative reciprocal" namely \(-\frac{1}{4}\) this will require working with fractions

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

would it (-1/4)(2) + x = x -5/2

OpenStudy (anonymous):

you are missing a y somewhere

OpenStudy (anonymous):

point is \((2,-3)\) use \(y-y_1=m(x-x_1)\) with \(x_1=2,y_1=-3, m=-\frac{1}{4}\)

OpenStudy (anonymous):

y - (-3) = -1/4(x - 2)

OpenStudy (anonymous):

yes, that is the correct first step

OpenStudy (anonymous):

although you might want to go to \(y+3=-\frac{1}{4}(x-2)\) or else make that the second step

OpenStudy (anonymous):

now distribute the \(-\frac{1}{4}\) on the right

OpenStudy (anonymous):

y + 11/4 = (x - 2

OpenStudy (anonymous):

just work on the right hand side, not the left

OpenStudy (anonymous):

so would it be x = -7/4

OpenStudy (anonymous):

\[y+3=-\frac{1}{4}(x-2)\] \[y+3=-\frac{1}{4}x+(-\frac{1}{4})\times (-2)\]by the distributive property then \[y+3=-\frac{1}{4}x+\frac{1}{2}\]

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

then subtract \(3\) from both sides to get \(y\) by itself you get \[y=-\frac{1}{4}x+\frac{1}{2}-3\] or \[y=-\frac{1}{4}x-\frac{5}{2}\]

OpenStudy (anonymous):

don't forget you need an \(x\) and a \(y\) in your answer your answer needs to look like \(y=mx+b\) this is not a "solve for \(x\)" problem, you do not want to end up with \(x=\text{some number}\)

OpenStudy (anonymous):

will I do anything to it, if so am I adding 5/2 to each side

OpenStudy (anonymous):

no it is good like it is it looks just like \(y=mx+b\) where \(m=-\frac{1}{4}\) and \(b=-\frac{5}{2}\)

OpenStudy (anonymous):

oooooooooooooh hold on

OpenStudy (anonymous):

oh okay I got what your saying now

OpenStudy (anonymous):

i am sorry, it says in "standard form"

OpenStudy (anonymous):

lets start here \[y+3=-\frac{1}{4}(x-2)\] and make it look like \(ax+by=c\)

OpenStudy (anonymous):

multiply both sides by 4 to clear the fraction. you get \[4y+12=-(x-2)\] or \[4y+12=-x+2\] add \(x\) to both sides, gives \[x+4y+12=2\] subtract \(12\) form both sides you get \[x+4y=-10\] which looks just like \(ax+by=c\)

OpenStudy (anonymous):

thank you this was less omplicated than the other lol